Diffusion Morning Report 03/29/11

Ted Berg tberg at stanford.edu
Tue Mar 29 06:28:52 PDT 2011


Good Morning All,
     Looking good this morning. Thermco Poly1 is running after 
resolution of growth rate issues. Epi is running.TEOS pump is due back 
today. Thermco LTO needs a test after repairs. Thermco 2 had no TLC 
flow. Ebeam compressor is dead. No other issues at this time.

					
	Date: 	3/29/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	*has been run* 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*pump should be back today* 	*.25/30* 	** 	*1*
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*20.67/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*9.6/30* 	** 	**
Tylan 11-LTO 	*yellow* 	** 	*3.28/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*yellow* 	*needs test after repair* 	** 	** 	**
AMAT Epi 	*yellow* 	*running* 	** 	** 	**
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 600 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110329/c9990d93/attachment.html>


More information about the diffusion mailing list